# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

### 18 Dec 2023 v2

@@ 5.1. The Effectiveness of Local-Aware Knowledge (RQ1)
Local Information can assist LLMs in code generation by making them aware of local contextual information. We extract and organize four types of Local-Aware Knowledge from the local module: local functions, class instance attributes, local module FQNs, and module variables. These four types of essential knowledge provide different perspectives on the context of the local module. We aim to investigate whether Local-Aware Information can genuinely help LLMs reuse local functions during code generation. Additionally, we also aim to identify the optimal local knowledge configuration, determining whether all four categories of knowledge should be made available to the model’s awareness or only specific ones.
@@
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in Table 1.
@@
Table 1. Comparison of the results obtained with different Local Information. And find the best configuration for Local Information Extraction.
@@
ChatGPT 0 0 0 0.754
ChatGPT + Local Functions 0.5 0.559 0.528 0.780
ChatGPT + Local Functions              +Class Instance Attributes 0.563 0.623 0.592 0.809
ChatGPT + Local Functions              +Class Instance Attributes              +Module Variables 0.489 0.567 0.525 0.783
ChatGPT +Local Functions              +Class Instance Attributes              +Module Variables              +Local Module FQN 0.542 0.629 0.582 0.809
5.1.1. Result Analysis
Table 1 displays the results corresponding to code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. vanilla ChatGPT has a precision and recall of 0 because it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
When we provide the model with knowledge of Local Functions, it immediately gains the ability to utilize local functions, resulting in a Precision of 0.558, a Recall of 0.624, an F1 of 0.525, and an accuracy of 0.779. This is because knowledge of Local Functions enables the model to directly invoke local functions, reducing code redundancy and enhancing code reusability and efficiency.
Upon additionally providing knowledge of Class Instance Attributes, the model’s Precision, Recall, and F1 score improve by 12.78%, 11.44%, and 12.15%, respectively, compared to providing only knowledge of Local Functions. This improvement is attributed to the provision of Class Instance Attributes, which enables the model to better understand how local functions interact with class variables in the project, leading to more accurate application of local functions.
However, when we introduce the Module Variables, the model’s performance drastically declines, particularly with a 13.21% drop in Precision. This may be due to the increased information load resulting from knowledge of Module Variables, which negatively impacts performance.
Nevertheless, the model’s performance is restored and approaches the capabilities of providing knowledge of Local Functions and Class Instance Attributes when we subsequently add knowledge of Local Module FQN. This suggests the existence of subtle interactions within the knowledge configuration. When the model receives only one or two key knowledge types, its performance may be superior to when it receives more information. However, as the volume of knowledge increases to a certain extent, the model’s capabilities strengthen, resulting in improved performance.
Ultimately, we opt for the Local-Aware Knowledge configuration, which includes knowledge of Local Functions and Class Instance Attributes. This configuration will be used as the default in the following experiments.
@@ 5.1.1. Result Analysis
5.2. The effectiveness of Global-Module-Aware Knowledge (RQ2)
In this RQ, we aim to investigate whether global-aware information can genuinely help LLMs reuse local functions during code generation. Additionally, we seek to determine the optimal top-k value for retrieving candidate functions, obtaining the best global-aware configuration.
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total ive different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in section 5.1, where each result is evaluated for whether it uses global functions.
Table 2. Comparison of the results obtained for different Top k numbers of retrieved functions to find the best k value. Where Precision, Recall, and F1 are for global functions.
@@ 5.2. The effectiveness of Global-Module-Aware Knowledge (RQ2)
No Global-Aware (k = 0) 0 0 0 0 0.819
k=1 1.757 0.535 0.441 0.484 0.788
k=5 8.154 0.658 0.533 0.589 0.825
k=10 16.031 0.550 0.441 0.490 0.793
k=15 23.592 0.521 0.418 0.464 0.783
5.2.1. Result Analysis
Table 2 presents the results obtained using ”No Global-Aware” and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.535 and 0.442, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 16.76%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ 5.2.1. Result Analysis
5.3. The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In section 3.2.3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
Table 3. Compare the difference between having Third-Party-Library Aware and not having it on Precision, Recall, F1, Library Coverage.
Third-Party-Library Awareness Precision Recall F1 Accuracy Library Coverage
No 0.464 0.620 0.530 0.676 0.686
Yes 0.599 0.901 0.719 0.815 0.906
5.3.1. Result Analysis
Table 3 demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 29.14%, 45.44%, 35.66%, and 20.47%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 32.09%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinations [59, 60, 61].
Finding 3: Third-Party-Library Aware information significantly improves the model’s ability to utilize available third-party libraries, reducing the risk of potential compatibility issue.
5.4. The Effectiveness of $A^{3}$odGen Framework (RQ4)
In previous sections, we have individually verified the effectiveness of Local-Aware, Global-Aware, and Third-Party-Library-Aware information. We also explored the optimal configurations. In this section, we aim to integrally validate the effectiveness of our $A^{3}$-CodGen Framework from various perspectives. We compare our $A^{3}$-CodGen with three baselines: vanilla ChatGPT, Local-Aware Generation, and $A^{2}$ Generation (Local-Aware and Global-Aware). In addition to metrics based on the confusion matrix, we also calculate library coverage, line of code (LOC), and CodeBLEU. Among these metrics, a lower LOC indicates simpler code, and a higher CodeBLEU suggests that the generated code closely resembles the ground truth.
Table 4. Comparison of the overall effect of Framework with Ablation Baselines
@@ 5.4. The Effectiveness of $A^{3}$odGen Framework (RQ4)
Local Reuse Precision 0 0.564 0.719 0.701
Recall 0 0.624 0.570 0.654
F1 0 0.592 0.636 0.677
Accuracy 0.755 0.810 0.794 0.830
Global Reuse Precision 0 0 0.658 0.648
Recall 0 0 0.533 0.605
F1 0 0 0.589 0.626
Accuracy 0.820 0.820 0.825 0.856
Third-Party Library Reuse Precision 0.8 0.687 0.464 0.599
Recall 0.492 0.563 0.620 0.901
F1 0.609 0.619 0.530 0.719
Accuracy 0.598 0.668 0.676 0.815
Library Coverage 0.792 0.855 0.686 0.905
Avg Line of Code (LOC) 16.320 16.487 12.598 12.804
CodeBLEU 0.339 0.344 0.312 0.322
5.4.1. Result Analysis
As shown in Table 4, our $A^{3}$-CodGen performs exceptionally well in terms of Local Reuse, Global Reuse, and Third-Party-Library Reuse, demonstrating its capability to effectively leverage information from the code repository. Furthermore, the framework generates code with optimal line of codes (LOC), indicating its capability to produce concise and efficient code.
Regarding CodeBLEU, we observed an interesting phenomenon: as our model’s code reuse capability increases, the value decreases. After manual inspection, we found that the unwanted excessive code reuse led to the decline of CodeBLEU. Specifically, as our model tends to make full use of existing code and functions by design, it generates distinctly different but more concise code compared to ground truth. Based on our manual evaluation, we surprisingly found that a large amount of generated code is better than the ground truth, i.e., our generated code subtly reuses existing functions, making the code more efficient, readable, and maintainable. For example, the function named allcap_differential_getter determines whether the input span contains mixed-case words. In the ground truth, a longer code snippet is written to accomplish this task (10 lines), but our approach generates a shorter, more streamlined code snippet that achieves the same goal by invoking a global function src.asent.getters.__get_lower_txt in another code file. This further underscores the effectiveness of our framework, which not only generates code similar to human developers but also assists developers in recommending more reusable code snippets.
@@ 5.4.1. Result Analysis
Finding 4: $A^{3}$-CodeGen framework excels in awaring three types of code repository knowledge, resulting in an excellent performance in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding reuse capabilities, our framework can generate concise code that aligns with the code repository’s working environment. $A^{3}$-CodeGen maximizes code reuse, and to some extent, the quality of the generated code surpasses that of the ground truth.
6. Discussion
@@
New Paradigm for Code Generation Tasks. Previously, code generation tasks aimed to automate code production to meet specific requirements, and the advent of LLMs like GitHub CoPilot and ChatGPT has substantially mastered this task. However, practical code generation tasks in real-world scenarios extend beyond mere requirement satisfaction. Generated code is expected to be aligned with project-specific requirements and leverage existing knowledge simultaneously, leading to better readability, maintainability, and practicability, which can further help developers in the software development process. Therefore, we hope this work sheds a light on this emerging but significant code generation topic: generating context-aware code that is compatible with the working environment.
@@
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality results [62, 63, 64]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In Section 5.1.1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in section 5.4, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This highlights the interplay between different types of code knowledge and emphasizes the importance of effectively mining and leveraging code knowledge from libraries in future research.
Threats to Validity. The threats to validity of this paper lies in three aspects: 1) Subjective Labeling. The data used in our study involves manual labeling, which introduces subjectivity in the annotations. We take a rigorous approach by enlisting the expertise of two professional developers to label the code. Furthermore, any instances with ambiguous labels are discussed and resolved to establish a reliable ground truth.
2) Limited Dataset Selection. We conduct experiments using a selection of 29 Python code repositories. While this number may appear limited, it’s essential to note that the repositories are carefully chosen based on their high quality, as indicated by a substantial number of stars and forks. These repositories are considered representative of the Python programming ecosystem. Additionally, to ensure the credibility of our ground truth, we employ rigorous sampling techniques to maintain a confidence level of 95% or higher. 3) General Applicability of Our Framework. Our framework is generic in the sense that its design is based on generalized code generation principles and natural language processing techniques, and does not depend on the specific nature of a particular programming language or domain. Although we chose Python as an example in our study, this was done only to validate the effectiveness of the framework.
7. Related Work
@@
Repository-Level Code Completion. Compared to traditional function-level code completion, repository-level code completion has drawn attention to practical developments [65, 38, 36, 66]. Shrivastava et al. generated example-specific prompts using a prompt proposal classifier [65] and further proposed RepoFusion to incorporate relevant repository context by training code models [38]. In addition, Ding et al. [36] proposed a cross-file context finder tool, named CCFinder, which constructs a project context graph and retrieves the most relevant cross-file context based on local import statements. RepoCoder [66] retrieved similar code snippets from the repository as part of a prompt and proposes an iterative retrieval-generation paradigm. This novel strategy bridged the gap between the retrieved context and the intended target through an interactive process.
Retrieval Augmented Code Generation. Retrieval-Augmented Code Generation (RACG) [67, 64, 68, 69] represents an emerging and promising area of research that harnesses the strengths of retrieval-based methods in conjunction with code generation techniques. Hayati et al. [64] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al. [68] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al. [67] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al. [69] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al. [70] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@
8. Conclusion
Code generation within code repositories often necessitates comprehensive knowledge of the entire repository context. Existing techniques often lack such a holistic understanding of code repository knowledge, which is crucial for effective code generation. In this paper, we introduce the AAA Code Generation framework, which is designed to construct, mine, and unify three essential types of code repository knowledge: local-aware, global-aware, and third-party-library aware knowledge. To validate the effectiveness of these knowledge types, we conducted benchmark experiments and identified optimal configurations for their utilization. Our experiments demonstrate that our framework significantly enhances the ability of Language Model Models (LLMs) to fully utilize local, global, and third-party library information. This results in the generation of concise, high-quality code. In the future, we plan to expand the applicability of our framework to multiple programming languages. Furthermore, we aim to enhance our knowledge utilization methods to enable LLMs to become more aware of various types of code knowledge within code repositories.

### 03 Mar 2024 v3

@@ V-A The Effectiveness of Local-Aware Knowledge (RQ1)
Local information aids LLMs in code generation by offering awareness of local contextual details. We extract and organize four types of Local-Aware Knowledge from the local module: local functions, class instance attributes, local module FQNs, and module variables. These knowledge types offer diverse perspectives on the context of the local module. Our goal is to assess whether local-aware knowledge genuinely enhances LLMs’ ability to reuse local functions in code generation. Additionally, we aim to identify the optimal configuration, determining whether all four knowledge categories should be available to the model’s awareness or only specific ones.
@@
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in Table I.
@@
TABLE I: Comparison of the Local Reuse Awareness results obtained with different Local Information. And find the best configuration for Local Information Extraction.
@@
Vallina ChatGPT N/A N/A N/A 0.715
ChatGPT +Local Functions 0.667 0.514 0.580 0.789
ChatGPT +Local Functions +Class Instance Attributes 0.742 0.633 0.683 0.833
ChatGPT +Local Functions +Class Instance Attributes +Module Variables 0.704 0.523 0.600 0.802
ChatGPT +Local Functions +Class Instance Attributes +Module Variables +Local Module FQN 0.747 0.541 0.628 0.817
V-A1 Result Analysis
Table I shows the reuse awareness results for code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. Vanilla ChatGPT exhibits a precision and recall of 0 as it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
When given knowledge of Local Functions, the model gains the ability to utilize them, resulting in a Precision of 0.667, Recall of 0.514, F1 score of 0.580, and accuracy of 0.789. This is because knowledge of Local Functions enables the model to directly invoke local functions. Upon additionally providing knowledge of Class Instance Attributes, the model’s Precision, Recall, and F1 score improve by 11.24%, 23.15%, and 17.76%, respectively, compared to providing only knowledge of Local Functions. This improvement is attributed to the provision of Class Instance Attributes, which enables the model to better understand how local functions interact with class variables in the project, leading to more accurate application of local functions.
However, when we introduce the Module Variables, the model’s performance drastically declines, particularly with a 17.38% drop in Recall. This may be due to the increased information load resulting from knowledge of Module Variables, which impacts performance. Nevertheless, the model’s performance is restored and approaches the capabilities of providing knowledge of Local Functions and Class Instance Attributes when we subsequently add knowledge of Local Module FQN. This suggests the existence of subtle interactions within the knowledge configuration. When the model receives only one or two key knowledge types, its performance may be superior to when it receives more information. However, as the volume of knowledge increases to a certain extent, the model’s capabilities strengthen, resulting in improved performance.
@@ V-A1 Result Analysis
V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
In this RQ, our goal is to explore the effectiveness of global-aware information in aiding LLMs to reuse global functions in code generation. We aim to identify the optimal top-k value for retrieving candidate functions to achieve the best global-aware configuration.
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total five different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in section V-A, where each result is evaluated for whether it uses global functions.
TABLE II: Comparison of the Global Reuse Awareness results obtained for different Top k numbers of retrieved functions to find the best k value.
@@ V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
No Global-Aware (k = 0) N/A N/A N/A N/A 0.833
k=1 1.757 0.384 0.516 0.440 0.781
k=5 8.154 0.542 0.703 0.612 0.851
k=10 16.031 0.493 0.562 0.526 0.830
k=15 23.592 0.395 0.531 0.453 0.786
V-B1 Result Analysis
Table II presents the results obtained using “No Global-Aware” and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.384 and 0.516, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 14.05%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ V-B1 Result Analysis
V-C The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In section III-B3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
TABLE III: Compare the difference between having Third-Party-Library Aware and not having it on Third-Party Library Reuse Awareness and Library Coverage.
3rd-Party-Lib Awareness Precision Recall F1 Accuracy Library Coverage
No 0.727 0.427 0.538 0.713 0.875
Yes 0.957 0.587 0.727 0.828 0.940
V-C1 Result Analysis
Table III demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 31.64%, 37.47%, 35.13%, and 16.13%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 7.43%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinations [57, 58, 59].
Finding 3: Third-Party-Library Aware information significantly improves the model’s ability to utilize available third-party libraries, reducing risk of potential compatibility issue.
V-D The Effectiveness of $A^{3}$odGen Framework (RQ4)
In previous sections, we have individually verified the effectiveness of Local-Aware, Global-Aware, and Third-Party-Library-Aware information. We also explored the optimal configurations. In this section, we aim to integrally validate the effectiveness of our $A^{3}$-CodGen Framework from various perspectives. We compare our $A^{3}$-CodGen with three baselines: vanilla ChatGPT, $A^{1}$ Generation (Local-Aware), and $A^{2}$ Generation (Local-Aware and Global-Aware). We evaluate our framework using four metrics: reuse awareness, reuse correctness, library coverage, and lines of code (LOC). The F1 score serves as a comprehensive measure of reuse awareness. Notably, a lower LOC signifies more concise code.
TABLE IV: Comparison of the overall effect of Framework with Ablation Baselines
@@ V-D The Effectiveness of $A^{3}$odGen Framework (RQ4)
Local Reuse Awareness (F1) N/A 0.683 0.664 0.693
Correctness N/A 0.648 0.646 0.633
Global Reuse Awareness (F1) N/A N/A 0.612 0.632
Correctness N/A N/A 0.604 0.601
3rd-Party Library Reuse Awareness (F1) 0.640 0.659 0.538 0.727
Correctness 0.640 0.742 0.743 0.850
Library Coverage 0.786 0.863 0.875 0.940
Avg Line of Code (LOC) 16.308 16.507 12.598 12.804
V-D1 Result Analysis
Upon evaluating the awareness aspect, as shown in Table IV, $A^{3}$-CodGen demonstrates excellence in local reuse, global reuse, and third-party library reuse. This underscores the effective assistance provided by the three categories of repository-aware knowledge in enabling the model to be cognizant of repository information during code generation.
In terms of correctness, we observe that Local Reuse Correctness is optimal for $A^{1}$ Generation (Local-Aware), Global Reuse Correctness is best for $A^{2}$ Generation (Local-Aware and Global-Aware), and Third-Party Library Reuse Correctness is highest for $A^{3}$-CodGen. This suggests that increasing input knowledge may influence the model’s performance in specific reuse scenarios. Nevertheless, $A^{3}$-CodGen exhibits only a slight difference from the baseline in local reuse and global reuse correctness (0.003-0.015). Despite this minor difference in correctness metrics and its superior performance in other aspects, $A^{3}$-CodGen demonstrates robust overall performance in local reuse, global reuse, and third-party library reuse.
Looking at the Library Coverage, $A^{3}$-CodGen achieves near-perfect performance with a score of 0.940, signifying its ability to effectively reuse pre-installed libraries and avoid compatibility issues such as “ModuleNotFoundError: No module named xxx.” Furthermore, the framework consistently generates code with an optimal number of lines (LOC), indicating its capability to produce concise and efficient code.
@@ V-D1 Result Analysis
Finding 4: $A^{3}$-CodeGen framework excels in awaring three types of code repository knowledge, resulting in an excellent performance in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding reuse capabilities, our framework can generate concise code that aligns with the code repository’s working environment.
VI Discussion
@@
New Paradigm for Code Generation Tasks. Code generation tasks have traditionally focused on automating code production for specific requirements, with advancements like GitHub CoPilot and ChatGPT mastering this task. However, real-world code generation extends beyond meeting requirements; it requires generating code aligned with project-specific needs and leveraging existing knowledge for improved readability, maintainability, and practicability. Therefore, we hope this work sheds a light on this emerging but significant code generation topic: generating context-aware code that is compatible with the working environment.
@@
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality results [60, 61, 62]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In Section V-A1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in section V-D, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This underscores the interaction of diverse code knowledge types and stresses the significance of efficiently mining and utilizing code knowledge in future studies.
Threats to Validity. The threats to validity of this paper lies in two aspects: 1) Subjective Labeling. Our study involves manual labeling, introducing subjectivity in annotations. We mitigate this by engaging two professional developers for labeling, resolving any instances with ambiguous labels to establish a reliable ground truth. The dataset is public for community validation [19]. 2) Limited Dataset Selection. Experiments use 29 meticulously selected Python code repositories, emphasizing high quality indicated by stars and forks, while avoiding data leaks. These repositories collectively represent the Python programming ecosystem. Rigorous sampling ensures a confidence level of 95% or higher, ensuring the credibility of the ground truth. In the future, we’ll maintain this dataset, providing the community with a valuable resource for evaluating repository-aware code generation.
VII Related Work
@@
Repository-Level Code Completion. Repository-level code completion is gaining traction over traditional function-level completion [63, 39, 37, 64]. Shrivastava et al. generated example-specific prompts using a prompt proposal classifier [63] and further proposed RepoFusion to incorporate relevant repository context by training code models [39]. In addition, Ding et al. [37] proposed a cross-file context finder tool, named CCFinder, which constructs a project context graph and retrieves the most relevant cross-file context based on local import statements. RepoCoder [64] retrieved similar code snippets from the repository as part of a prompt and proposes an iterative retrieval-generation paradigm. This novel strategy bridged the gap between the retrieved context and the intended target through an interactive process.
Retrieval Augmented Code Generation. Retrieval Augmented Code Generation (RACG) [65, 62, 66, 67] represents an emerging and promising area of research that combines the strengths of retrieval-based methods with code generation techniques. Hayati et al. [62] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al. [66] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al. [65] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al. [67] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al. [68] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@
VIII Conclusion and Future Work
Code generation within code repositories often necessitates comprehensive knowledge of the entire repository context. Existing techniques often lack such a holistic understanding of code repository knowledge, which is crucial for effective code generation. In this paper, we introduce the $A^{3}$-CodGen framework, which is designed to construct, mine, and unify three essential types of code repository knowledge: local-aware, global-aware, and third-party-library-aware knowledge. To validate the effectiveness of these knowledge types, we conducted benchmark experiments and identified optimal configurations for their utilization. The results demonstrate that the $A^{3}$-CodGen framework significantly enhances the ability of Language Model Models (LLMs) to fully utilize local, global, and third-party library information. This results in the generation of concise, high-quality code. In the future, we plan to expand the applicability of our framework to multiple programming languages. Furthermore, we aim to enhance our knowledge utilization methods to enable LLMs to become more aware of various types of code knowledge within code repositories.